OpenStudy (mathmath333):

In 2 bags there are to be put altogether 5 red & 12 white balls,neither bag being empty. How must the balls be divided so as to give a person who draws one ball from either bag the least chance of drawing a red ball ?

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ In 2 bags there are to be put altogether 5 red & 12 white }\hspace{.33em}\\~\\ & \normalsize \text{ balls,neither bag being empty. How must the balls be }\hspace{.33em}\\~\\ & \normalsize \text{ divided so as to give a person who draws one ball from}\hspace{.33em}\\~\\ & \normalsize \text{ either bag the least chance of drawing a red ball ?}\hspace{.33em}\\~\\ & a.)\ \dfrac{3}{35} \hspace{.33em}\\~\\ & b.)\ \dfrac{5}{32} \hspace{.33em}\\~\\ & c.)\ \dfrac{7}{32} \hspace{.33em}\\~\\ & d.)\ \dfrac{1}{16} \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (michele_laino):

I'm very sorry, I don't know your answer

OpenStudy (mathmath333):

answer given is 5/32 ,

OpenStudy (anonymous):

Ya

OpenStudy (anonymous):

Yes

OpenStudy (error1603):

Try B

OpenStudy (anonymous):

B

OpenStudy (amistre64):

this is the closest thing i can come up with on the google

OpenStudy (mathmate):

We could reason as follows: There are 5 red and 12 white. We have to put at least one ball in each bag. Say we put a red in one, and nothing in the other, then the probability of drawing a red would be: (1/1+4/16)/=0.75. Can we do better? If we exchange and put a white in bag A, and the rest in the other (B), then probability =(0+5/16)/2 = 0.15625 At this point, moving any white ball to bag A will increase the probability of drawing in B, so is moving any red. So the distribution A(0,1) and B(5,11) is a local minimum giving P(R)=5/32. I have not yet found a way to show that this is also the global minimum without brute force (which is what I did). Hope someone would give it a try.

OpenStudy (mathmath333):

in book it is given , \(\large \color{black}{\begin{align} & \dfrac12 \times \dfrac{5}{16}+ \dfrac12 \times 0= \dfrac{5}{32} \hspace{.33em}\\~\\ \end{align}}\) which i didnt understand

OpenStudy (mathmate):

Yes, that's the same as : (0+5/16)/2 = 0.15625 = 5/32

OpenStudy (mathmate):